Parth (parthkohli):

OwlKid's puzzle. A yard sale is selling Lay's packets each for $1. For every three waste wrappers you give him, you get another packet. How many Lay's packets can you buy if you have $15?

OpenStudy (anonymous):

I didn't know you liked rap music Parth ;-) (typo?) 1 packet per $1 = rate 3 packet's wrappers = +1 more packet (for free, yes)? \[f(d) = \lfloor \frac{4}{3}\cdot (1 \frac{packet}{$})\cdot (d \ \ $) \rfloor\] I think? I've never been very strong with these types of problems

OpenStudy (anonymous):

That's a floor function btw, truncates, since you can't have a third of a packet. Find f(15)

OpenStudy (anonymous):

Also Parth, I thought you hated integrals?

Parth (parthkohli):

Wow. I prefer it manually.

Parth (parthkohli):

And yeah, for free.

OpenStudy (unklerhaukus):

twenty one?

Parth (parthkohli):

Close, but not correct.

OpenStudy (unklerhaukus):

25

Parth (parthkohli):

Hmm. Still close.

OpenStudy (unklerhaukus):

hmmm

OpenStudy (anonymous):

22...

Parth (parthkohli):

Right there. What water said.

OpenStudy (unklerhaukus):

$15=15p+15w 15w=5p+5w 3w=1p 1p+2w=1p $15=15p+5p+1p+1p=22p

Parth (parthkohli):

\( \color{Black}{\Rightarrow $15 \Longrightarrow 15packs \Longrightarrow 15 wrappers \Longrightarrow +5packs \Longrightarrow 3 + 2 wrappers}\) \(\Longrightarrow +1pack + 2wrappers \Longrightarrow 1 wrapper + 2 wrappers \Longrightarrow 3 \)wrappers. 15 + 5 + 1 + 1 = 22

OpenStudy (anonymous):

See, we bought 15 packets with $15 directly.. Now out of 15 packets we have 15 wrappers.. So, for every 3 wrapper we get 1 so 15 wrappers will give us 5 more packets.. Now, we have 5 wrappers, out of 5 wrappers, we will get 1 more packet.. Now, this 1 packet and that remaining two wrapper will make 3 wrappers that is 1 more packet.. So, 15 + 5 + 1 + 1 = 22

Parth (parthkohli):

Great job, both of you!